Shift report 20000725 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Tue Jul 25 2000 - 08:08:43 EDT

  • Next message: http account: "Shift report 20000725 08:00-16:00"

    ------------------------------------------------------------
    Date: Tue Jul 25 08:08:42 2000, Shift: 00:00-08:00
    Supervisor: PC, with: PS
    Log book#3 pages 70 - 75
    
    Shift Summary:
    ------------------------------------------------------------
    Beam most of the night @ gamma = 70
    run 1927 126 trigger 6 / 1100
    run 1928 876 trigger 6 of 50k 
    run 1929 1345 / 140k @ 7:45 - still running
    Intensity has been high in both beams.
    
    Tracks have been seen in all TPCs!!!! Someone with a nice 
    program should take a look at T2 data. 2 pictures can be found 
    in ~operator/t2pics. It would perhaps also be interesting to
    compare with data where we did not have anode voltage on when
    doing pedestals and tonight where we had anode voltage on.
    2 TPC HVs tripped when they lost the beam in run 1928.
    ZDC ntuples looked like old ntuples with 1 neutron peak.
    
    Taked to Konstantin about DAQ errors in T3 map when we arrived.
    He stressed that it was important to include all detectors 
    camac, fb1, fb3, fb4, mtpc1, mtpc2, tpc1, tpc2 in
    pedestal runs. Should the TPCs also be in for CAMAC pedestal
    runs ?
    
    With beam 2 nights in a row maybe it is time to turn on the
    magnets ?
    
    Problems :
    Supermonitor crashed with a storage exhausted. Anybody using 
    new in their event method should take a look.
    Printer did not want to print anything from operator after
    ~3 am.
    Paper for the printer is needed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Tue Jul 25 2000 - 08:12:15 EDT